OpenStudy (amtran_bus):

A stone is thrown horizontally at 20 m/s from the top of a cliff 63 m high. How fast is it moving the instant before it hits the ground? a. 33 m/s b. 40 m/s c. 38 m/s d. 29 m/s

OpenStudy (amtran_bus):

Help please

OpenStudy (amtran_bus):

They said use A^2+B^2???

OpenStudy (anonymous):

You know that Ek = Ep: mv^2 / 2 = mgh Now solve for the vertical velocity v. g is the gravitational constant 9.8 and h is the height in meters

OpenStudy (anonymous):

let A = v (the vertical velocity you get) and B = 20 (horizontal velocity) Then the total velocity is: \[\sqrt{A^{2}+B^{2}}\]